# Dynamic planning in hierarchical active inference

### v2

@@ 4.1 Dynamic inference by model reduction
This kind of dynamic inference has several utilities, e.g., it can be used to infer which one among multiple objects an agent is following – as exemplified in Figure 14 – by generating trajectories for different objects and comparing them with the one it is perceiving [69]. As a last note, the dynamics precisions of Equations 33 and 38 have here an interesting interpretation specular to the observation precisions $\bm{\pi}_{o}$. Active inference and predictive coding assume that whenever an agent perceives high noise about a sensory modality, the precision of that generative model will decrease because it cannot be trusted for understanding the state of affairs of the world [11, 12]. In addition, the dualism between action and perception inherent to the free energy principle tells us that the optimization of precisions – which are thought to be encoded as synaptic gains – could play a crucial role in attention mechanisms that selectively sample sensory data [105, 106]. Based on this assumption, we could interpret a low precision $\bm{\pi}_{x,m}$ as the decreased agent’s confidence over that intention for minimizing the prediction errors in the current context; however, a low intention precision could also mean that the agent does not intend to rely on it for the realization of a desired goal. In short, there is a dual interpretation of intention precisions about explaining a situation (e.g., the result of a grasping action to understand an object far away from the hand) or solving a task (e.g., grasping an object when it is out of reach). This perspective unveils an additional mechanism besides the fast inference of hidden causes that we mentioned before: a slow learning of reduced precisions that lets the agent score – and, crucially, focus on – those intentions that would be appropriate for a specific scenario [69].
@@ 4.2 A discrete interface for dynamic planning
Numerous studies have demonstrated that the brains of athletes are marked by a higher activation of posterior and subcortical regions that involves little or no conscious thinking, producing fluid transitions between different motions; in contrast, the brain of a novice requires a higher demand of prefrontal computations that results in lower performances [107, 108, 109]. From an active inference perspective, we can compare the proficiency of athletes with the continuous model of Figure 5(b), corresponding to the subcortical sensorimotor loops. This model encodes a transition mechanism that is not very flexible, but which precisely for this reason allows it to react much more rapidly to environmental stimuli, e.g., when grasping objects moving at high speed [75]. In general, this strategy can be very effective when the environment has limited uncertainty and the task to be solved comprises a rigid sequence of actions, which the agent has already correctly learned. However, suppose that the agent is introduced to a novel task, or to a highly complex task that requires careful thinking about the imminent future. In this case, it should be capable of replanning the correct sequence of actions if something goes unexpected, and a high-level belief always producing an a-priori-determined behavior for the hidden causes would fall short in completing the task.
Figure 15: Interface between a discrete model (at the top) and several hybrid units (at the bottom). For clarity, the hidden states factorization of each unit is not displayed. The discrete hidden states $\bm{s}$ at time $\tau+1$ are computed – from the current hidden states $\bm{s}_{\tau}$ – by choosing some policy $\bm{\pi}_{\pi}$ related to a specific transition distribution encoded in $\bm{B}$. The best policy at any moment is the sequence of actions that is most likely to minimize the free energy $\mathcal{G}$ that the agent expects to perceive in the future. As a result, the agent will try to sample those observations that conform to its preference $\bm{C}$. The discrete hidden causes $\bm{v}^{(i)}$ at time $\tau$ are directly generated from discrete hidden states $\bm{s}_{\tau}$ through likelihood matrices $\bm{A}^{(i)}$, thus affording dynamic planning, synchronized behavior, and inference with multiple evidences.
Having replaced the continuous hidden causes of Figure 5(b) with discrete hidden causes in Figure 13(b), we can now endow the agent with planning capabilities through a discrete model composed of the following distributions – as shown in Figure 15:
$p(\bm{s}_{1:T},\bm{v}_{1:T},\bm{\pi})=p(\bm{s}_{1})p(\bm{\pi})\prod_{\tau}p(\bm{v}_{\tau}|\bm{s}_{\tau})p(\bm{s}_{\tau}|\bm{s}_{\tau-1},\bm{\pi})$ (40)
@@
Here, $\bm{A}$, $\bm{B}$, $\bm{D}$ are the likelihood matrix, transition matrix, and prior, $\bm{\pi}$ are the policies – which are not state-action mappings as in RL but sequences of actions – with prior $\bm{E}$, and $\bm{s}_{\tau}$ are the discrete hidden states at time $\tau$. These quantities have strict analogies with their continuous counterparts of Equation 3, i.e., the likelihood function $\bm{g}$, the dynamics function $\bm{f}$, the prior $\bm{\eta}_{x}$, and the hidden causes $\bm{v}$, with the difference that the hidden states do not encode instantaneous paths expressed in generalized coordinates, but sequences of future states defined by discrete variables. Here, we wish to infer the posterior distribution:
$p(\bm{s}_{1:T},\bm{\pi}|\bm{v}_{1:T})=\frac{p(\bm{v}_{1:T}|\bm{s}_{1:T},\bm{\pi})p(\bm{s}_{1:T},\bm{\pi})}{p(\bm{v}_{1:T})}$ (42)
@@
As before, this requires computing the intractable model evidence $p(\bm{v}_{1:T})$, so we resort to a variational approach: expressing the approximate posterior by its sufficient statistics $\bm{s}_{\pi,\tau}$ and conditioning upon a specific policy:
$\begin{split}p(\bm{s}_{1:T}|\bm{v}_{1:T},\bm{\pi})&\approx q(\bm{s}_{1:T},\bm{\pi})=q(\bm{\pi})\prod_{\tau}^{T}q(\bm{s}_{\tau}|\bm{\pi})\\ q(\bm{\pi})&=Cat(\bm{\pi})\\ q(\bm{s}_{\tau}|\bm{\pi})&=Cat(\bm{s}_{\pi,\tau})\end{split}$ (43)
@@
we infer the most likely discrete hidden states at time $\tau$ by computing the gradient of the related free energy $\mathcal{F}_{\pi}$ of that policy:
$\bm{s}_{\pi,\tau}=\sigma(\ln\bm{B}_{\pi,\tau-1}\bm{s}_{\pi,\tau-1}+\bm{B}_{\pi,\tau+1}^{T}\bm{s}_{\pi,\tau+1}+\sum_{i}\ln\bm{A}^{(i)^{T}}\bm{v}^{(i)}_{\tau})$ (44)
where we applied a softmax function to ensure that it is a proper probability distribution. Here, $\bm{A}^{(i)}$ is the likelihood matrix that sends predictions for the $i$th hybrid unit. In fact, if we connect several units to the discrete model, each of them has an independent interface whereby the discrete model computes different signals and waits for the next step $\tau+1$, when it can infer its hidden states based on multiple accumulated evidences. Recall that in the combined structure of Figure 15, the role of the hybrid unit was to predict a dynamic trajectory from a discrete intention prior, and to infer the most likely intention in a continuous period $T$. But now the intention prior is generated from a high-level policy that decides which action to take next in the current situation:
$\bm{v}_{\tau}=\sigma(\ln\bm{A}\bm{s}_{\tau}+\bm{l}_{\tau})$ (45)
where $\bm{l}_{\tau}$ is the bottom-up message at time $\tau$. The inference of the policies additionally considers unobserved outcomes as random variables, finding the most likely sequence of actions that will lead to some preferred outcomes. More formally, the policy posterior $q(\bm{\pi})$ is found by comparing the policy prior with the expected free energy $\mathcal{G}$, which is defined as the free energy that the agent expects to perceive in the future:
$\bm{\pi}=\sigma(\ln\bm{E}-\mathcal{G})$ (46)
Assuming that an agent has some preference $\bm{C}$ about future outcomes, the expected free energy $\mathcal{G}_{\pi}$ under policy $\pi$ will consist of a pragmatic or goal-seeking term toward that preference, and an epistemic or uncertainty-reducing term (see [55] for more details):
$\begin{split}\mathcal{G}_{\pi}&\approx\sum_{\tau}D_{KL}[q(\bm{v}_{\tau}|\bm{\pi})||p(\bm{v}_{\tau}|\bm{C})]+\operatorname*{\mathbb{E}}_{q(\bm{s}_{\tau}|\bm{s}_{\tau-1},\bm{\pi})}[H[p(\bm{v}_{\tau}|\bm{s}_{\tau})]]\\ &=\sum_{\tau}\bm{v}_{\pi,\tau}(\ln\bm{v}_{\pi,\tau}-\bm{C}_{\tau})+\bm{s}_{\pi,\tau}\bm{H}_{A}\end{split}$ (47)
@@
Note in the above equations that the likelihood matrix $\bm{A}$ expresses a conditional probability over the discrete hidden causes $\bm{v}_{\tau}$. As in conventional hybrid models, the discrete hidden states are linked to the hidden causes, but now the latter directly act as discrete observations generated by the likelihood matrix, which thus replaces the prior $\bm{H}_{v}$ in Equation 31. In sum, computing the posterior probability over policies $\bm{\pi}$ turns into finding the best action that makes the agent conform to the dual objective defined by $\mathcal{G}$. Here, the discrete actions are not intended as actual motor commands similar to Equation 9, but as abstract actions over high-level representations. In fact, the hierarchical nature of discrete models in active inference makes it possible to perform decision-making with a separation of temporal scales, wherein a specific level can generate and infer the states and the paths of the level below [110, 111, 112]. Further evaluating the consequences of an action for a longer time horizon affords more advanced planning called sophisticated inference [113]. Computing actions with the free energy of the future is different from the motor control of continuous models, which only minimize the free energy of present states.
@@
Figure 16: In this task, the agent (a 4-DoF arm with an additional 4-DoF hand composed of two fingers) has to pick a moving object, and place it at a home position. The agent has a shallow structure with a single IE module computing the hand position from the arm joint angles. Note that the object belief is rapidly inferred, and as soon as the picking action is complete, the belief is gradually pulled toward the home position, resulting in a second reaching movement. The top right panel shows the hand-object distance over time, while the bottom right panel displays the dynamics of the discrete action probabilities used to infer the next discrete state. See [75] for more details.
@@
In addition to the previous agents, it is now possible to synchronize the behavior of different continuous signals based on the same high-level policy. For instance, one can realize a pick-and-place operation with a moving object – as represented in Figure 16 – producing smooth transitions between reaching and grasping actions, respectively performed in extrinsic and intrinsic domains [75]. Note that an intermediate phase between the two actions naturally arises, corresponding to a composite approaching movement. In principle, the learning of intention precisions $\bm{\pi}_{x,m}$ (not to be confused with the policy notation) might shed light on how motor skill learning occurs, via message passing between continuous intentions and discrete policies. Moreover, through this kind of dynamic planning the agent can infer and realize instantaneous trajectories even within the same discrete period $\tau$, useful, e.g., for grasping the moving object without waiting for the successive replanning step. Finally note that to correctly maintain a goal state, we now need to introduce a hidden cause loosely corresponding to the stay action commonly used in discrete tasks [56]. This hidden cause can be linked to an identity intention, i.e., $\bm{i}_{stay}(\bm{x})=\bm{x}$, which can be interpreted as the agent’s desire to maintain the current state of affairs of the world [69]. Again, the dualism between action and perception also relates this hidden cause to the initial stationary state of the task, and translates into a specular desire for a phase of pure perceptual inference – as the one shown in the simulation of Figure 6.
4.3 Deep hybrid models
@@
Figure 17 portrays a deep hybrid model designed for solving a flexible tool-use task [99]. It combines the expressivity of a (deep) hierarchical formulation, the advantages of inferring and imposing dynamic multi-step intentions inherent to a hybrid unit, and the possibility of encoding external objects and other agents. As in Figure 15, the IE modules communicate with a discrete model at the top, but now they are combined in a hierarchical fashion recapitulating the agent’s kinematic chain. As a consequence, two different goal-directed strategies arise. Considering a simple reaching movement, an attractor imposed at the hand level would generate a cascade of extrinsic prediction errors flowing back to the previous levels and finding a suitable kinematic configuration with the hand over the target. This corresponds to a horizontal hierarchical depth occurring along the hybrid units, and can be compared to the process of motor babbling typical of infants [114], whereby random attractors are generated at different hierarchical levels to identificate the correct body structure. In addition to this naive strategy, since a discrete model can now generate intentions for every IE module (in both intrinsic and extrinsic domains), a more advanced behavior can be achieved once inverse kinematics is correctly performed, which imposes a specific path to the whole kinematic chain. This corresponds to a vertical hierarchical depth with two (discrete and continuous) temporal scales, steering the lower-level inferential process in a direction that, e.g., avoids singularities or gets out from local minima generated by repulsive attractors.
@@
There is thus a delicate balance between forward and backward extrinsic likelihood, and the top-down modulation of the discrete model:
$\dot{\bm{\mu}}_{e}^{(i)}\propto-\bm{\pi}_{e}^{(i-1)}\bm{\varepsilon}_{e}^{(i-1)}+\partial\bm{g}_{e}^{T}\bm{\pi}_{e}^{(i)}\bm{\varepsilon}_{e}^{(i)}+\partial\bm{\eta}^{\prime(i)T}_{x,e}\bm{\pi}_{x,e}^{(i)}\bm{\varepsilon}_{x,e}^{(i)}$ (49)
where $\partial\bm{\eta}^{\prime(i)T}_{x,e}$ is the gradient of the trajectory prior of Equation 36. From the discrete model’s perspective, the discrete hidden states produce a specific combination of hidden causes for each hybrid unit; this combination generates a composite trajectory in the continuous domain weighting independent intentions, taking into account dynamic elements for the whole discrete step. After this period, evidence is accumulated for every hybrid unit, eventually inferring the most likely discrete state that may have generated the trajectories of the self and the environment.
A non-trivial issue exists in tasks requiring tool use, e.g., reaching a ball with the extremity of a stick. Much as other agents may have different kinematic structures than the self, a tool may have its own hierarchy (e.g., even a simple stick is represented by two Cartesian positions and an angle) that must somehow be integrated into the agent’s generative model. Specifically, reaching an object with the extremity of a tool means defining a potential kinematic chain augmented by a new virtual level, letting the agent think of the tool as an extension of its arm. This is possible by linking the two visual observations of the tool to the hand and virtual levels in a second pathway of the hidden states, as shown in Figure 17. Since the intrinsic units of the IE modules also encode information about limb lengths, the agent can infer through visual observations not only its kinematic structure, but also the actual length of the tool [115]. While this second pathway is still marked by a clear distinction between the tool and the arm since the hand level receives observations from both elements, a third pathway is constructed such that the observation of the ball is only linked to the virtual level. As a result, this new potential configuration views the arm and the tool as being part of the same kinematic chain. The interactions between these three pathways (shown in Figure 18) may shed light on how the remapping of the motor cortex gradually occurs with extensive tool use [6, 7], modifying the boundaries between the self and the environment.

### v3

@@ 4.1 Dynamic inference by model reduction
As a last note, the dynamics precisions of Equations 33 and 38 have here an interesting interpretation equivalent to the observation precisions $\bm{\Pi}_{o}$. Predictive coding assumes that whenever an agent perceives high noise about a sensory modality, the precision of that generative model will decrease because it cannot be trusted for understanding the state of affairs of the world [19, 20]. In addition, the dualism between action and perception inherent to the free energy principle tells us that the optimization of precisions – which are thought to be encoded as synaptic gains – could play a crucial role in attention mechanisms that selectively sample sensory data [118, 119]. Based on these assumptions, we note a dual interpretation of (reduced) dynamics precisions. A low precision $\bm{\Pi}_{x,m}$ related to a grasping action could mean that it is unreliable to explain the current context (e.g., an object far away from the hand). In addition, it could mean that the agent does not intend to rely on it to achieve a goal (e.g., grasping an object when it is out of reach). This perspective unveils an additional mechanism besides the fast inference of hidden causes that we mentioned before: a slow learning of reduced precisions that lets the agent score – and, crucially, focus on – those intentions that would be appropriate for a specific scenario [65].
@@ 4.2 A discrete interface for dynamic planning
Numerous studies have demonstrated that the brains of athletes are marked by a higher activation of posterior and subcortical regions that involves little or no conscious thinking, producing fluid transitions between different motions; in contrast, the brain of a novice requires a higher demand of prefrontal computations that results in lower performances [120, 121, 122]. From an active inference perspective, we can compare the proficiency of athletes with the continuous model of Figure 5(a), corresponding to the subcortical sensorimotor loops. This model encodes a transition mechanism that is not very flexible, but that precisely for this can react much more rapidly to environmental stimuli, e.g., when grasping objects moving at high speed [66]. In general, this strategy can be very effective when the environment has limited uncertainty and the task to be solved comprises a rigid sequence of actions, which the agent has already correctly learned. However, suppose that the agent is introduced to a novel or complex task that requires careful thinking about the imminent future. In this case, it should be capable of replanning the correct sequence of actions if something goes unexpectedly, and a high-level belief always producing an a-priori-determined behavior for the hidden causes would fail to complete the task.
Figure 15: Interface between a discrete model (at the top) and several hybrid units (at the bottom). For clarity, the hidden states factorization of each unit is not displayed. The discrete hidden states $\bm{s}$ at time $\tau+1$ are computed from the current hidden states $\bm{s}_{\tau}$, by choosing some policy $\bm{\pi}_{\pi}$ related to a specific transition distribution encoded in $\bm{B}$. The best policy at any moment is the sequence of actions that most minimizes the expected free energy $\mathcal{G}$. As a result, the agent will try to sample those observations that conform to its preference $\bm{C}$. The discrete hidden causes $\bm{v}^{(i)}$ at time $\tau$ are directly generated from discrete hidden states $\bm{s}_{\tau}$ through likelihood matrices $\bm{A}^{(i)}$, thus affording dynamic planning, synchronized behavior, and inference with multiple evidences.
Having replaced the continuous hidden causes of Figure 5(a) with discrete hidden causes in Figure 13(c), we can now endow the agent with planning capabilities through a discrete model composed of the following distributions – as shown in Figure 15:
$p(\bm{s}_{1:\mathcal{T}},\bm{v}_{1:\mathcal{T}},\bm{\pi})=p(\bm{s}_{1})p(\bm{\pi})\prod_{\tau}^{\mathcal{T}}p(\bm{v}_{\tau}|\bm{s}_{\tau})p(\bm{s}_{\tau}|\bm{s}_{\tau-1},\bm{\pi})$ (40)
@@
Here, $\bm{A}$, $\bm{B}$, $\bm{D}$ are the likelihood matrix, transition matrix, and prior of the hidden states, $\bm{\pi}$ are the policies (which are not state-action mappings as in RL but sequences of actions), $\bm{s}_{\tau}$ are the discrete hidden states at time $\tau$. These quantities have strict analogies with their continuous counterparts of Figure 1(b), i.e., the likelihood function $\bm{g}$, the dynamics function $\bm{f}$, the prior $\bm{\eta}_{x}$, hidden causes $\bm{v}$, and hidden states $\tilde{\bm{x}}$, with the difference that the discrete hidden states do not encode instantaneous paths expressed in generalized coordinates, but sequences of future states. Crucially, $\mathcal{G}$ is the expected free energy, defined as the free energy that the agent expects to perceive in the future.
Note here that the likelihood matrix $\bm{A}$ expresses a conditional probability over the discrete hidden causes $\bm{v}_{\tau}$. As in conventional hybrid models, the discrete hidden states are linked to the hidden causes, but now the latter directly act as discrete outcomes generated by the likelihood matrix, which thus replaces the prior $\bm{H}_{v}$ in Equation 31. Here, we wish to infer the posterior distribution:
$p(\bm{s}_{1:\mathcal{T}},\bm{\pi}|\bm{v}_{1:\mathcal{T}})=\frac{p(\bm{v}_{1:\mathcal{T}}|\bm{s}_{1:\mathcal{T}},\bm{\pi})p(\bm{s}_{1:\mathcal{T}},\bm{\pi})}{p(\bm{v}_{1:\mathcal{T}})}$ (42)
@@
As before, this requires computing the intractable model evidence $p(\bm{v}_{1:\mathcal{T}})$, so we resort to a variational approach: we first express the approximate posterior by its sufficient statistics $\bm{s}_{\pi,\tau}$ and conditioning upon a specific policy:
$p(\bm{s}_{1:\mathcal{T}}|\bm{v}_{1:\mathcal{T}},\bm{\pi})\approx q(\bm{s}_{1:\mathcal{T}},\bm{\pi})=q(\bm{\pi})\prod_{\tau}^{\mathcal{T}}q(\bm{s}_{\tau}|\bm{\pi})$ (43)
where:
$\begin{split}q(\bm{\pi})&=Cat(\bm{\pi})\\ q(\bm{s}_{\tau}|\bm{\pi})&=Cat(\bm{s}_{\pi,\tau})\end{split}$ (44)
We then compute the variational free energy $\mathcal{F}_{\pi}$ of that policy:
$\mathcal{F}_{\pi}=\sum_{\tau=1}^{\mathcal{T}}\bm{s}_{\pi,\tau}\ln\bm{s}_{\pi,\tau}-\sum_{\tau=1}^{\mathcal{T}}\bm{v}_{\tau}\ln\bm{A}\bm{s}_{\pi,\tau}-\bm{s}_{\pi,1}\ln\bm{D}-\sum_{\tau=2}^{\mathcal{T}}\bm{s}_{\pi,\tau}\ln\bm{B}_{\pi,\tau}\bm{s}_{\pi,\tau-1}$ (45)
@@
Differentiating Equation 45 allows us to infer the most likely discrete hidden states at time $\tau$ for policy $\pi$:
$\bm{s}_{\pi,\tau}=\sigma(\ln\bm{B}_{\pi,\tau-1}\bm{s}_{\pi,\tau-1}+\ln\bm{B}_{\pi,\tau}^{T}\bm{s}_{\pi,\tau+1}+\ln\bm{A}^{T}\bm{v}_{\tau})$ (46)
where we applied a softmax function to ensure that it is a proper probability distribution. In order to infer the policies, we additionally consider unobserved outcomes as random variables, and then condition the joint probability of the generative model upon some preferences $\bm{C}$. In this way, the agent can predict future outcomes and select the most likely sequence of actions that will lead it to its desired state. The expected free energy $\mathcal{G}_{\pi}$ under policy $\pi$ consists of a pragmatic or goal-seeking term toward the agent’s preferences, and an epistemic or uncertainty-reducing term (see [96] for more details):
$\begin{split}\mathcal{G}_{\pi}&\approx\sum_{\tau}^{\mathcal{T}}D_{KL}[q(\bm{v}_{\tau}|\bm{\pi})||p(\bm{v}_{\tau}|\bm{C})]+\operatorname*{\mathbb{E}}_{q(\bm{s}_{\tau}|\bm{s}_{\tau-1},\bm{\pi})}[H[p(\bm{v}_{\tau}|\bm{s}_{\tau})]]\\ &=\sum_{\tau}^{\mathcal{T}}\bm{v}_{\pi,\tau}(\ln\bm{v}_{\pi,\tau}-\bm{C}_{\tau})+\bm{s}_{\pi,\tau}\bm{H}_{A}\end{split}$ (47)
@@
Here, $\bm{v}_{\pi,\tau}$ is the prediction of the hidden causes at time $\tau$ and under policy $\pi$, $\bm{C}_{\tau}$ is the logarithm of the prior preference over the hidden causes, and $\bm{H}_{A}$ is the ambiguity of the likelihood matrix $\bm{A}$. The policy $\bm{\pi}$ is then found by applying a softmax function to the expected free energy, as defined in Equation 41.
Using the (deep) hierarchical model outlined in the previous section, we may connect several hybrid units to the discrete model by appropriate likelihood matrices $\bm{A}^{(i)}$. Each of them has an independent interface whereby the discrete model computes different signals and waits for the next step $\tau+1$, when it can infer its hidden states based on multiple accumulated evidences. Recall that in the combined structure of Figure 15, the role of the hybrid unit was to predict a trajectory from a discrete intention prior, and to infer the most likely intention in a continuous period $T$. But now the intention prior is generated from the discrete model:
$\bm{v}_{\tau}=\sigma(\ln\bm{A}\bm{s}_{\tau}+\bm{l}_{\tau})$ (49)
where $\bm{l}_{\tau}$ is the bottom-up message at time $\tau$ found by Equation 38, and we compute $\bm{s}_{\tau}$ by marginalizing over all policies, i.e. $\bm{s}_{\tau}=\sum_{\pi}\pi_{\pi}\bm{s}_{\pi,\tau}$.
@@
In sum, computing the posterior probability over policies $\bm{\pi}$ turns into finding the best action that conforms to the dual objective defined by $\mathcal{G}$. Here, the discrete actions are not intended as actual motor commands similar to Equation 7, but as abstract actions over high-level representations. In fact, the hierarchical nature of discrete models in active inference permits performing decision-making with a separation of temporal scales, wherein a specific level can generate and infer the states and the paths of the level below [123, 124, 125]. Further evaluating the consequences of an action for a longer time horizon affords more advanced planning called sophisticated inference [126]. Computing actions with the expected free energy is different from the motor control of continuous models, which only minimize the free energy of present states.
@@
Figure 16: In this task, the agent (a 4-DoF arm with an additional 4-DoF hand composed of two fingers) has to pick a moving object, and place it at a goal position. The agent has a shallow structure with a single IE module computing the hand position from the arm joint angles. Besides the spatial dimension, the agent can easily integrate multisensory information such as touch, dimensions of objects, and other more abstract properties. Note that the object belief is rapidly inferred, and as soon as the picking action is complete, the belief is gradually pulled toward the goal position, resulting in a second reaching movement. The top right panel shows the hand-object distance over time, while the bottom right panel displays the dynamics of the discrete action probabilities used to infer the next discrete state. See [66] for more details.
@@
In addition to the previous agents, it is now possible to synchronize the behavior of different continuous signals based on the same high-level policy. For instance, one can realize a pick-and-place operation with a moving object – as represented in Figure 16 – producing smooth transitions between reaching and grasping actions, respectively performed in extrinsic and intrinsic domains [66]. Note how an intermediate phase between the two actions naturally arises, corresponding to a composite approaching movement. In principle, the learning of dynamics precisions $\bm{\Pi}_{x,m}$ might shed light on how motor skill learning occurs, via message passing between continuous trajectories and discrete policies. Moreover, through this kind of dynamic planning the agent can infer and realize instantaneous trajectories even for the same continuous period $T$ within a discrete step $\tau$, useful, e.g., for grasping the moving object without waiting for the successive replanning step. Third, this configuration permits learning a discrete representation of the environment based on continuous evidence: as shown in [127], learning the likelihood matrix $\bm{A}$ involves counting the coincidences between targets and body movements, thus adapting a response strategy (risky versus conservative) depending on the difficulty of the context. Additionally, through learning of the prior $\bm{D}$, a habitual behavior toward the chosen decision emerges. Similarly, one could update preferred states encoded in the matrix $\bm{C}$ based on current observations and actions.
Since the agent plans with trajectories and not positions, to correctly maintain a goal state we need to introduce a hidden cause loosely corresponding to the stay action commonly used in discrete tasks [97]. This hidden cause can be linked to an identity map, i.e., $\bm{i}_{stay}(\bm{x})=\bm{x}$, which can be interpreted as the agent’s desire to maintain the current state of the world [65]. This hidden cause also relates to the initial state of the task, and translates into a phase of pure perceptual inference.
4.3 Flexible hierarchies
@@
Figure 17 portrays a deep hybrid model designed for solving a tool-use task [90]. It combines the expressivity of a (deep) hierarchical formulation, the advantages of planning trajectories inherent to a hybrid unit, and the possibility of encoding object affordances and other agents. As in Figure 15, the IE modules communicate with a discrete model at the top, but now they are combined in a hierarchical fashion recapitulating the agent’s kinematic chain. As a consequence, two different goal-directed strategies arise. Considering a simple reaching movement, an attractor imposed at the hand level would generate a cascade of extrinsic prediction errors flowing back to the previous levels and finding a suitable kinematic configuration with the hand over the target. This corresponds to a horizontal hierarchical depth occurring along the hybrid units, and can be compared to the process of motor babbling typical of infants [128], whereby random attractors are generated at different hierarchical levels to identify the correct body structure. In addition to this naive strategy, since a discrete model can now generate trajectories for every IE module (in both intrinsic and extrinsic domains), a more advanced behavior can be achieved once inverse kinematics is correctly performed, which imposes a specific path to the whole kinematic chain. This corresponds to a vertical hierarchical depth with two (discrete and continuous) temporal scales, steering the lower-level inference in a direction that, e.g., avoids singularities or gets out from local minima generated by repulsive attractors.
@@
Considering the update rule for the (extrinsic) hybrid unit at level $i$:
$\dot{\bm{\mu}}_{x,e}^{(i)}\propto-\bm{\Pi}_{o,e}^{(i-1)}\bm{\varepsilon}_{o,e}^{(i-1)}+\partial\bm{g}_{e}^{T}\bm{\Pi}_{o,e}^{(i)}\bm{\varepsilon}_{o,e}^{(i)}+\partial\bm{\eta}^{\prime(i)T}_{x,e}\bm{\Pi}_{x,e}^{(i)}\bm{\varepsilon}_{x,e}^{(i)}$ (50)
where $\partial\bm{\eta}^{\prime(i)T}_{x,e}$ is the gradient of the trajectory prior of Equation 36, we note a delicate balance between forward and backward extrinsic likelihood, and the top-down modulation of the discrete model. From the discrete model’s perspective, the discrete hidden states produce a specific combination of hidden causes for each hybrid unit; this combination generates a composite trajectory in the continuous domain weighting distinct potential trajectories, taking into account dynamic elements for the whole discrete step $\tau$. After this period, evidence is accumulated for every hybrid unit, eventually inferring the most likely discrete state that may have generated the actual trajectory, related to the self and the environment.
A non-trivial issue exists in tasks requiring tool use, e.g., reaching a ball with the extremity of a stick. Much as other agents may have different kinematic structures than the self, a tool may have its own hierarchy (e.g., even a simple stick is represented by two Cartesian positions and an angle) that must somehow be integrated into the agent’s generative model. Specifically, reaching an object with the extremity of a tool means defining a potential body configuration augmented by a new virtual level. This new level does not exist in the generative process and the belief of the actual body configuration $\bm{x}_{0}$. Nonetheless, if we consider the two potential configurations, the agent thinks of the tool as an extension of its arm, hence flexibly modifying its body schema as discussed in the Introduction. This is possible by linking the two visual observations of the tool to the hand and virtual levels in a second pathway of the hidden states, as shown in Figure 17. Since the intrinsic units of the IE modules also encode information about limb lengths, the agent can not only infer its kinematic structure through visual observations, but also the actual length of the tool [129]. While this second pathway is still marked by a clear distinction between the tool and the arm since the hand level receives observations from both elements, a third pathway is constructed such that the observation of the ball is only linked to the virtual level. As a result, this new potential configuration views the arm and the tool as part of the same kinematic chain. The interactions between these three pathways (shown in Figure 18) may shed light on how the remapping of the motor cortex gradually occurs with extensive tool use [10, 11], modifying the boundaries between the self and the environment.